OpenStudy (anonymous):

Can a tangent with the slope -(1/2) be drawn on the curve y=log4x ( that's a base 4). Explain, with calculus.

OpenStudy (anonymous):

yes, set f'(x) = -1/2

OpenStudy (anonymous):

Sorry but could you further explain

OpenStudy (anonymous):

do you know how to find derivative of log_4(x) ?

OpenStudy (anonymous):

not really, is it 1/xln4?

OpenStudy (anonymous):

very good. set 1/[x ln(4) ] = -1/2 and solve for x

OpenStudy (anonymous):

wait why do you set them equal to each other?

OpenStudy (anonymous):

because you're try to find the where on the graph is the slope -1/2

OpenStudy (anonymous):

m okay I got x=-2/ln4 is that correct?